# Sliding DFT-based Signal Recovery for Modulo ADC with 1-bit Folding Information

### v3

@@ III-F Comparison with Conventional ADCs
To demonstrate the advantage of a modulo ADC with 1-bit folding side information over a conventional ADC, we first derive the MSE of a conventional $b$-bit ADC under the non-subtractive dithered quantization framework. We compare the modulo ADC and conventional ADC under the same bit budget used for amplitude quantization. For a triangle dither $d[n]\in\left(-\frac{2\lambda}{2^{b}},+\frac{2\lambda}{2^{b}}\right)$, the ADC parameter $\lambda$ should be set to $\lambda=\left(1+\frac{1}{2^{b}}\right)\|f(t)\|_{\infty}$. After digital filtering, the quantization noise power (which is also the MSE) becomes
@@
Comparing the MSE guarantee for modulo ADC in Theorem 1 and the derived MSE for a conventional ADC, it can be observed that $\mathrm{MSE}_{\mathrm{mod}}=\mathcal{O}\left(\frac{1}{\mathrm{OF}^{2}}\right)$ while $\mathrm{MSE}_{\mathrm{conv}}=\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$ under sufficiently large $b$. The fast decay rate of $\mathrm{MSE}_{\mathrm{mod}}$ with respect to $\mathrm{OF}$ is due to the reduction of the ADC parameter $\lambda$. Smaller ADC range entails smaller quantization bins. This demonstrates the superior performance of modulo ADCs over conventional ADCs in oversampled systems. Still, the modulo ADC has a slightly higher hardware complexity due to the folding operation and the 1-bit side information $c[n]$.
@@ III-G Computational Complexity
We now derive the computational complexity of the proposed algorithm in terms of the signal length $N_{0}$, window length $N$, window roll-off $\alpha$ ($0\leq\alpha\leq 1$) and oversampling factor $\mathrm{OF}=\frac{1}{\rho}$. The number of segments depends on the ratio $\frac{N_{0}}{N}$, and the parameter $\alpha$ dictates how large is the overlap between consecutive segments. For large $N_{0}$, the number of segments to be processed by the algorithm is in $\mathcal{O}\left(\frac{N_{0}}{N\left(1-\alpha/2\right)}\right)$. Looking at the complexity of the residual recovery algorithm at each frame, the calculation of $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\in\mathbb{C}^{|\mathcal{S}_{i}|\times K}$ is the most computationally-expensive operation. More precisely, this calculation is in $\mathcal{O}\left(K|\mathcal{S}_{\max}|^{2}\right)$, where $|\mathcal{S}_{\max}|=\max_{i}|\mathcal{S}_{i}|$. Since $K\approx(1-\rho)N$ and $|\mathcal{S}_{\max}|\leq K$ under appropriate choice of $\mathrm{OF}$ and $\lambda^{\prime}$, then the complexity of computing $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}$ is in $\mathcal{O}\left((1-\rho)^{3}N^{3}\right)$. Considering all the segments to be processed, the overall computational complexity of the proposed recovery method is in $\mathcal{O}\left(\frac{N_{0}(1-\rho)^{3}N^{2}}{1-\alpha/2}\right)$. Hence, the computational complexity of our proposed algorithm is linear in the signal length $N_{0}$.
Based on the derived computational complexity of the proposed algorithm, the speed of the proposed sliding DFT-based recovery scheme can be improved by choosing a small oversampling factor (high $\rho$) that satisfies the sufficient condition in Theorem 1. However, reducing $\mathrm{OF}$ negatively impacts the MSE as shown in equation (17). The speed of the proposed sliding DFT-based recovery scheme can also be improved by choosing a small window length $N$. However, a small window length also needs a higher $\alpha$ to compensate for the increased spectral leakage. Finally, we note that the recovery algorithm in [14, 48] is a special case of our proposed recovery algorithm with $N=N_{0}$ and $\alpha=0$, i.e., rectangular window with no overlap. Hence, its complexity is in $\mathcal{O}((1-\rho)^{3}N_{0}^{3})$. Since $N$ is typically chosen to be much smaller than $N_{0}$, our proposed sliding DFT-based recovery is $\mathcal{O}\left(\frac{N_{0}^{2}(1-\alpha/2)}{N^{2}}\right)$ faster than the recovery algorithm in [14]. Suppose the signal to be processed has $N_{0}=100,000$ samples. Our proposed algorithm with sliding window length $N=256$ and roll-off parameter $\alpha=0.25$ can unfold the whole signal at roughly $7.48\times 10^{-6}$ times the total time needed by the recovery method in [14, 48] to unfold the whole signal.
@@ IV Numerical Results
Figure 5: Empirical evaluation of $\log_{2}\left(1+0.75M\right)$ as a function of segment length (NN) and $|\mathcal{S}_{i}|$ for $\mathrm{OF}=4$.
Figure 6: Empirical evaluation of $\log_{2}\left(1+0.75M\right)$ as a function of segment length (NN) and $|\mathcal{S}_{i}|$ for $\mathrm{OF}=8$.
Figure 7: Empirical evaluation of $\log_{2}\left(1+0.75M\right)$ as a function of segment length (NN) and $|\mathcal{S}_{i}|$ for $\mathrm{OF}=12$.
In this section, we validate the MSE performance guarantees for the proposed algorithm established in the previous section by comparing the theoretical MSE results with the simulated performance. Numerical experiments are also conducted to gain additional insight on how the algorithm parameters affect the performance guarantees. We also compare the proposed sliding DFT recovery against the HoD-based recovery.
IV-A Analysis of the Sufficient Condition for $b$
We first investigate how the sufficient condition on $b$ (i.e., equation (16)) is affected by algorithm and input signal parameters such as segment length $N$, oversampling factor $\mathrm{OF}$, and number of non-zero elements of the 1-bit folding information signal $c[n]$. To this end, we consider three $\mathrm{OF}$ parameter settings ($\mathrm{OF}=4,\;8,\;12$) and three segment lengths (N=64, 128, 256N=64,\;128,\;256). We also consider three values of $|\mathcal{S}_{i}|$ ($|\mathcal{S}_{i}|=\frac{N}{32},\frac{N}{16},\frac{N}{8}$). For each combination of these parameter settings, we generate 100,000 realizations of $\mathbf{V}_{\mathcal{S}_{i}}$ by selecting $|\mathcal{S}_{i}|$ out of $N$ columns of $\mathbf{V}$. We then estimate the second term of equation (16) by calculating
@@ IV-A Analysis of the Sufficient Condition for $b$
Figures 5-7 depict the empirical evaluation $\log_{2}\left(1+0.75\tilde{M}\right) fo$r $\mathrm{OF}=4$, $\mathrm{OF}=8$, and $\mathrm{OF}=12$, respectively. This term accounts for the extra bit resolution needed by the proposed algorithm (in addition to the 3 bits in the first term of equation (16)) to satisfy the sufficient condition. Based on these evaluations, it can be seen that using a higher oversampling factor generally decreases the second term of equation (16). One insight that can be drawn from this observation is that a reduction in the modulo ADC resolution can be compensated by oversampling. Figures 5-7 also show that longer segment length and frequent crossing of the modulo thresholds increase $M$.
As a final note in this numerical experiment, the sufficient condition for $b$ derived in the previous section may not be tight. We show in the succeeding numerical experiment that the proposed algorithm can achieve the performance guarantees in Theorem 1 even if we set $b=4$.
IV-B Theoretical vs. Simulated MSE of the Sliding DFT-based Unfolding
@@
where $p_{\mathrm{rc}}(t)$ is a raised-cosine pulse with roll-off parameter $\beta=1$. For the numerical simulation, we set the filter span of the raised-cosine pulse to 20. The peak amplitude of the raised-cosine pulse $p_{\mathrm{rc}}(t)$ is set to unity. The pulse amplitude $A_{m}$ is drawn uniformly from the interval [0.5,+1][-0.5,+1], and $T=1$ second is the time difference between two consecutive pulses. Intuitively, the signal $f(t)$ is constructed by generating 50,000 raised-cosine pulses with random amplitudes at every symbol period $T$. Consequently, the maximum frequency is $\frac{\omega_{m}}{2}=\frac{\pi(1+\beta)}{T}$. The interval is chosen so that $f(t)$ has a non-zero mean value. We want to show that the algorithm can work on signals with non-zero average value. This signal is fed to a low-resolution 4-bit modulo ADC with sampling rate $\frac{1}{T_{\mathrm{s}}}=\frac{\mathrm{OF}\omega_{m}}{2\pi}$ to produce the quantized modulo samples $f_{\lambda^{\prime},\mathrm{q}}[n]$. The dither sequence and modulo threshold are configured according to the modulo ADC parameters.
@@
For the sliding DFT-based recovery algorithm, we set the roll-off parameter of the tapered cosine window to $\alpha=0.5$ and window length $N=64$. Note that the length of the tapered cosine window is also the segment length and the DFT size of the algorithm. To investigate the impact of spectral leakage, we consider two different spectral leakage values: $\delta_{\mathrm{SL}}=\frac{\pi}{32},\;\frac{\pi}{16}$ (resp. $K_{\mathrm{SL}}=4,8$). The spectral leakage in the discrete frequency interval $\left(\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}}\right)$ is assumed to be negligible. We will show later that this assumption is indeed valid for this simulation setup since the theoretical and simulated MSE performances coincide. Equation (18) is used to specify the modulo threshold $\lambda^{\prime}$. The simulated performance is evaluated by taking the squared difference between the true samples $f[n]$ and the signal unfolding output $\hat{f}[n]$ and averaging the results over all $N_{0}$ samples.
@@
Figure 8: Simulated and theoretical MSE vs. oversampling factor of the proposed sliding DFT-based unfolding algorithm for $N=64$, $\alpha=0.5$, $b=4$ and $K_{\mathrm{SL}}=4,8$. Superimposed in the figure are the simulated and theoretical MSE curves of a conventional ADC
@@
Figure 8 depicts the simulated MSE (in decibels) of the modulo ADC with sliding DFT recovery and its the theoretical MSE (based on Theorem 1) under the aforementioned settings. The oversampling factor is swept from $\mathrm{OF}=4$ to $\mathrm{OF}=50$. The numerical results for modulo ADC demonstrate that the derived MSE performance guarantee is accurate since it coincides with the simulated MSE for all $\mathrm{OF}$ and $K_{\mathrm{SL}}$ settings considered. It can be observed that larger $\delta_{\mathrm{SL}}$ increases the MSE of the sliding DFT unfolding method. This can be attributed to two things: (a) the widening of the passband region of the digital LPF used in the last step of the unfolding method, and (b) the increase in the $\lambda^{\prime}$. Consequently, higher quantization noise power would be present at the digital LPF output.
@@
The theoretical MSE curve for $K_{\mathrm{SL}}=0$ (i.e. spectral leakage neglected) is also superimposed in Figure 8. It is evident from the figure that the slope of the $K_{\mathrm{SL}}=0$ MSE curve is steeper at the high $\mathrm{OF}$ regime compared to those of $K_{\mathrm{SL}}=4$ and $K_{\mathrm{SL}}=8$ MSE curves. This validates the asymptotic MSE growth rates mentioned in Section III. Nonetheless, the modulo ADC which uses our sliding DFT recovery still outperforms the conventional ones even if there is spectral leakage. The gap between the MSE curves of the modulo ADC and conventional ADC widens as $\mathrm{OF}$ is increased. At $\mathrm{OF}=40$, the difference between the MSE incurred by the conventional ADC and the MSE incurred by the proposed sliding DFT-based recovery with $K_{\mathrm{SL}}=8$ is roughly 23 dB.
IV-C Comparison between the Sliding DFT-based recovery and HoD-based recovery

### v4

@@ III-F Comparison with Conventional ADCs
To demonstrate the advantage of a modulo ADC with 1-bit folding information over a conventional ADC, we first derive the MSE of a conventional $b$-bit ADC under a non-subtractive dithered quantization framework. We compared the modulo ADC and the conventional ADC under the same bit budget used for amplitude quantization. For a triangle dither $d[n]\in\left(-\frac{2\lambda}{2^{b}},+\frac{2\lambda}{2^{b}}\right)$, the ADC parameter $\lambda$ should be set to $\lambda=\left(1+\frac{1}{2^{b}}\right)\|f(t)\|_{\infty}$. After digital filtering, the quantization noise power (which is also the MSE) becomes
@@
Comparing the MSE guarantee for modulo ADC in Theorem 1 and the derived MSE for a conventional ADC, it can be observed that $\mathrm{MSE}_{\mathrm{mod}}=\mathcal{O}\left(\frac{1}{\mathrm{OF}^{2}}\right)$ while $\mathrm{MSE}_{\mathrm{conv}}=\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$ under sufficiently large $b$. The fast decay rate of $\mathrm{MSE}_{\mathrm{mod}}$ with respect to $\mathrm{OF}$ is due to the reduction of the ADC parameter $\lambda$. A smaller ADC range results in smaller quantization bins. This demonstrates the superior performance of modulo ADCs compared to conventional ADCs in oversampled systems. However, the modulo ADC has a slightly higher hardware complexity due to the folding operation and the 1-bit folding information $c[n]$. A detailed comparison of power consumption, cost, and temperature behavior requires specific circuit-level implementations of modulo ADCs and conventional ADCs, which is beyond the scope of this simulation-based study.
@@ III-G Computational Complexity
We now derive the computational complexity of the proposed algorithm in terms of signal length $N_{0}$, window length $N$, window roll-off $\alpha$ ($0\leq\alpha\leq 1$) and oversampling factor $\mathrm{OF}=\frac{1}{\rho}$. The number of segments depends on the ratio $\frac{N_{0}}{N}$, and parameter $\alpha$ dictates the extent of overlap between consecutive segments. For a large $N_{0}$, the number of segments to be processed by the algorithm is in $\mathcal{O}\left(\frac{N_{0}}{N\left(1-\alpha/2\right)}\right)$. Considering the complexity of the residual recovery algorithm in each frame, the calculation of $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\in\mathbb{C}^{|\mathcal{S}_{i}|\times K}$ is the most computationally-expensive operation. More precisely, this calculation is in $\mathcal{O}\left(K|\mathcal{S}_{\max}|^{2}\right)$, where $|\mathcal{S}_{\max}|=\max_{i}|\mathcal{S}_{i}|$. Since $K\approx(1-\rho)N$ and $|\mathcal{S}_{\max}|\leq K$ under appropriate choice of $\mathrm{OF}$ and $\lambda^{\prime}$, the complexity of computing $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}$ is in $\mathcal{O}\left((1-\rho)^{3}N^{3}\right)$. Considering all segments to be processed, the overall computational complexity of the proposed recovery method is in $\mathcal{O}\left(\frac{N_{0}(1-\rho)^{3}N^{2}}{1-\alpha/2}\right)$. Hence, the computational complexity of the proposed algorithm is linear in the signal length $N_{0}$.
Based on the derived computational complexity of the proposed algorithm, the speed of the proposed STFT-based recovery scheme can be improved by selecting a small oversampling factor (high $\rho$) that satisfies the sufficient condition in Theorem 1. However, reducing $\mathrm{OF}$ negatively impacts the MSE as shown in equation (22). The speed of the proposed STFT-based recovery scheme can also be improved by selecting a small window length $N$. However, a small window length also requires a higher $\alpha$ to compensate for the increased spectral leakage. Finally, we note that the recovery algorithm in [4, 5] is a special case of our proposed recovery algorithm with $N=N_{0}$ and $\alpha=0$, that is, a rectangular window with no overlap. Hence, its complexity is in $\mathcal{O}((1-\rho)^{3}N_{0}^{3})$. Because $N$ is typically chosen to be much smaller than $N_{0}$, our proposed STFT-based recovery is $\mathcal{O}\left(\frac{N_{0}^{2}(1-\alpha/2)}{N^{2}}\right)$ faster than the recovery algorithm in [4]. Suppose that the signal to be processed has $N_{0}=100,000$ samples. Our proposed algorithm with a sliding window length $N=256$ and roll-off parameter $\alpha=0.25$ can unfold the whole signal at roughly $7.48\times 10^{-6}$ times the total time needed by the recovery method in [4, 5] to unfold the entire signal.
@@ IV Numerical Results
In this section, we validate the MSE performance guarantees for the proposed algorithm established in the previous section by comparing the theoretical MSE results with the simulated performance. Numerical experiments are also conducted to gain additional insight on how the algorithm parameters affect the performance guarantees. We also compare the proposed STFT-based recovery method with the HoD-based recovery.
IV-A Analysis of the Sufficient Condition for ADC Resolution
We first investigate how the sufficient condition on $b$ (i.e., equation (21)) is affected by the algorithm and input signal parameters such as segment length $N$, oversampling factor $\mathrm{OF}$, and number of non-zero elements of the 1-bit folding information signal $c[n]$. To this end, we consider three $\mathrm{OF}$ parameter settings ($\mathrm{OF}=4,\;8,\;12$) and three segment lengths (N=64, 128, 256N=64,\;128,\;256). We also consider three values of $|\mathcal{S}_{i}|$ ($|\mathcal{S}_{i}|=\frac{N}{32},\frac{N}{16},\frac{N}{8}$). The selected window lengths and oversampling factor settings are chosen arbitrarily and are intended solely to demonstrate how these parameters affect the second term of the LHS of equation (21). For each combination of these parameter settings, we generate 100,000 realizations of $\mathbf{V}_{\mathcal{S}_{i}}$ by selecting $|\mathcal{S}_{i}|$ out of $N$ columns of $\mathbf{V}$. We then estimate the second term of equation (21) by calculating
@@ IV-A Analysis of the Sufficient Condition for ADC Resolution
Figure 7: Empirical evaluation of $\log_{2}\left(1+0.75M\right)$ as a function of segment length (NN) and $|\mathcal{S}_{i}|$ for $\mathrm{OF}=4$. The values of the segment length used in the numerical experiments are $N=64$, 128128, and $256$. The values of $|\mathcal{S}_{i}|/N$ are set to $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$.
Figure 8: Empirical evaluation of $\log_{2}\left(1+0.75M\right)$ as a function of segment length (NN) and $|\mathcal{S}_{i}|$ for $\mathrm{OF}=8$. The values of the segment length used in the numerical experiments are $N=64$, 128128, and $256$. The values of $|\mathcal{S}_{i}|/N$ are set to $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$.
Figure 9: Empirical evaluation of $\log_{2}\left(1+0.75M\right)$ as a function of segment length (NN) and $|\mathcal{S}_{i}|$ for $\mathrm{OF}=12$. The values of the segment length used in the numerical experiments are $N=64$, 128128, and $256$. The values of $|\mathcal{S}_{i}|/N$ are set to $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$.
Figures 7-9 depict the evaluation $\log_{2}\left(1+0.75\tilde{M}\right) fo$r $\mathrm{OF}=4,8$, and $12$, respectively. This term accounts for the extra bit resolution needed by the proposed algorithm (in addition to the 3 bits in the first term of equation (21)) to satisfy the sufficient condition. Based on these evaluations, it can be seen that using a higher oversampling factor generally decreases the second term of equation (21). One insight that can be drawn from this observation is that a reduction in the modulo ADC resolution can be compensated by oversampling. Figures 7-9 also show that longer segment lengths and frequent crossings of the modulo thresholds increase $M$.
To help interpret the observed trends, we recall that $M$ is the $\ell_{\infty}$-norm of the product of the Moore-Penrose inverse of $\mathbf{V}_{\mathcal{S}_{i}}$ and $\mathbf{V}_{\mathcal{S}_{i}^{c}}$. This quantity measures the maximum interference between the columns corresponding to folding indices and those outside the folding set. Therefore, the magnitude of $M$ is governed by two mechanisms: (i) the conditioning of $\mathbf{V}_{\mathcal{S}_{i}}$, which determines how much the pseudo-inverse operation amplifies the quantization noise, and (ii) the cross-coherence between $\mathbf{V}_{\mathcal{S}_{i}}$ and $\mathbf{V}_{\mathcal{S}_{i}^{c}}$. Increasing $\mathrm{OF}$ while keeping $|\mathcal{S}_{i}|$ fixed increases the dimension of the $|\mathcal{S}_{i}|$ column vectors of $\mathbf{V}_{\mathcal{S}_{i}}$. This leads to improved conditioning of $\mathbf{V}_{\mathcal{S}_{i}}$. In contrast, increasing $|\mathcal{S}_{i}|$ while keeping $\mathrm{OF}$ fixed has the opposite effect. The matrix $\mathbf{V}_{\mathcal{S}_{i}}$ becomes wider due to the increased number of column vectors, making $\mathbf{V}_{\mathcal{S}_{i}}$ more poorly conditioned. This results in a larger pseudo-inverse gain, which dominates the reduction in the number of columns of $\mathbf{V}_{\mathcal{S}_{i}^{c}}$.
As a final note in this numerical experiment, the sufficient condition for $b$ derived in the previous section may not be tight. In the succeeding numerical experiment, we show that the proposed algorithm can achieve the performance guarantees in Theorem 1, even if we set $b=4$.
IV-B Theoretical vs. Simulated MSE of the Proposed STFT-based Unfolding
@@
where $p_{\mathrm{rc}}(t)$ is a raised-cosine pulse with roll-off parameter $\beta=0.25$. For the numerical simulation, we set the filter span of the raised-cosine pulse to 20. The peak amplitude of the raised-cosine pulse $p_{\mathrm{rc}}(t)$ is set to unity. The pulse amplitude $A_{m}$ is drawn uniformly from the interval [0.5,+1][-0.5,+1], and $T=1$ second is the time difference between two consecutive pulses. Intuitively,signal $f(t)$ is constructed by generating 150,000 raised-cosine pulses with random amplitudes at every symbol period $T$. Consequently, the maximum frequency is $\frac{\omega_{m}}{2}=\frac{\pi(1+\beta)}{T}$. The interval is selected such that $f(t)$ has a non-zero mean value. We want to show that the algorithm can work on signals with a non-zero average value. This signal is fed to a low-resolution 4-bit modulo ADC with a sampling rate of $\frac{1}{T_{\mathrm{s}}}=\frac{\mathrm{OF}\omega_{m}}{2\pi}$ Hz to produce the quantized modulo samples $f_{\lambda^{\prime},\mathrm{q}}[n]$. The dither sequence and modulo threshold are configured according to the modulo ADC parameters.
@@
For the STFT-based recovery algorithm, we set the roll-off parameter of the tapered cosine window to $\alpha=0.50$ and window length $N=64$. This choice of window parameters provides a balance between computational complexity and spectral leakage suppression. Note that the length of the tapered cosine window is also the segment length and the DFT size of the algorithm. To investigate the impact of spectral leakage, we consider two different spectral leakage values: $\delta_{\mathrm{SL}}=\frac{3\pi}{64},\;\frac{4\pi}{64}$ (resp. $K_{\mathrm{SL}}=6,8$). Spectral leakage in the interval $\left(\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}}\right)$ is assumed to be negligible. We will show later that this assumption is valid for this simulation setup since the theoretical and simulated MSE performances coincide. Equation (23) is used to specify the modulo threshold $\lambda^{\prime}$. The simulated performance is evaluated by taking the squared difference between the true samples $f[n]$ and the signal unfolding output $\hat{f}[n]$ and averaging the results over all $N_{0}$ samples. Note that the simulated MSE is computed using a single realization of $f(t)$. This is sufficient because, when the conditions on $\mathrm{OF}$ and $b$ in Theorem 1 are satisfied, there is no unfolding error and the MSE is due to the quantization noise, which forms an ergodic process. Consequently, the statistical average is well-approximated by the time average of a single realization of $f(t)$ for a sufficiently large $N_{0}$.
@@
Figure 10: Simulated and theoretical MSE vs. oversampling factor of the proposed STFT-based unfolding algorithm for $N=64$, $\alpha=0.50$, $b=4$ and $K_{\mathrm{SL}}=6,8$. Superimposed in the figure are the simulated and theoretical MSE curves of a conventional ADC.
@@
Figure 10 depicts the simulated MSE (in decibels) of the modulo ADC with STFT-based recovery and its theoretical MSE (based on Theorem 1) under the aforementioned settings. The oversampling factor is swept from $\mathrm{OF}=4$ to $\mathrm{OF}=50$. The numerical results for the modulo ADC demonstrate that the derived MSE performance guarantee is accurate because it coincides with the simulated MSE for all $\mathrm{OF}$ and $K_{\mathrm{SL}}$ settings considered. It can be observed that larger $\delta_{\mathrm{SL}}$ increases the MSE of the STFT-based recovery method. This can be attributed to two factors: (a) the widening of the passband region of the digital LPF used in the last step of the unfolding method, and (b) the increase in $\lambda^{\prime}$ according to equation (23). Consequently, a higher quantization noise power would be present at the digital LPF output.
@@
The theoretical MSE curve for $K_{\mathrm{SL}}=0$ (i.e., spectral leakage neglected) is also superimposed in Figure 10. It is evident from the figure that the slope of the $K_{\mathrm{SL}}=0$ MSE curve is steeper in the high $\mathrm{OF}$ regime compared to those of $K_{\mathrm{SL}}=6$ and $K_{\mathrm{SL}}=8$ MSE curves. This validates the asymptotic MSE growth rates mentioned in Section III. Nonetheless, the modulo ADC, which uses our STFT-based recovery, still outperforms the conventional ADC, even in the presence of spectral leakage. The gap between the MSE curves of the modulo ADC and conventional ADC widened as the $\mathrm{OF}$ is increased. At $\mathrm{OF}=40$, the difference between the MSE incurred by the conventional ADC and that incurred by the proposed STFT-based recovery with $K_{\mathrm{SL}}=8$ is approximately 25 dB. This result demonstrates that modulo ADCs can outperform conventional ADCs in oversampled settings, potentially enabling lower power consumption by relaxing resolution requirements.
IV-C Comparison between the STFT-based recovery and HoD-based recovery